Required information

[The following information applies to the questions displayed below.]

Warnerwoods Company uses a perpetual inventory system. It entered into the following purchases and sales transactions for March.

Date Activities Units Acquired at Cost Units Sold at Retail
Mar. 1 Beginning inventory 240 units @ $53.80 per unit
Mar. 5 Purchase 295 units @ $58.80 per unit
Mar. 9 Sales 400 units @ $88.80 per unit
Mar. 18 Purchase 155 units @ $63.80 per unit
Mar. 25 Purchase 290 units @ $65.80 per unit
Mar. 29 Sales 270 units @ $98.80 per unit
Totals 980 units 670 units

3. Compute the cost assigned to ending inventory using (a) FIFO, (b) LIFO, (c) weighted average, and (d) specific identification. For specific identification, the March 9 sale consisted of 135 units from beginning inventory and 265 units from the March 5 purchase; the March 29 sale consisted of 115 units from the March 18 purchase and 155 units from the March 25 purchase.
Answer #1

A) FIFO

  

UNITS

RECIEPTS

UNIT COST

TOTAL COST ISSUES UNIT UNIT COSTS UNITS BALANCE UNIT COST TOTAL COST

OP BAL

MARCH 1

240 53.80 12912

MAR 5

295 58.80 17346

295

58.80 17346
535 30258
MAR 9

240

160

53.80

58.80

400 135 58.80

7938

MARCH 18 155 63.80 9889 155 63.80 9889
290 17827
MAR 25 290 65.80 19082 290 65.80 19082
580 36909
MARCH 29

135

135

58.80

63.80

20

290

63.80

65.80

1276

19082

20358

B) LIFO

  

MAR I

OP STOCK 240 @ 53.80   

MAR 5

PURCHASE 295 @58.80

MARCH 9

SALES 295 @ 58.80

105 @ 53.80 SO TOTAL 400

INVENTORY BALANCE

135 @ 53.80

MARCH 18

PURCHASE 155@ 63.80

MARCH 25

PURCHASE 290@ 65.80

INVENTORY BALANCE

   135 @ 53.80

155@ 63.80

   290@ 65.80

MARCH 29

SALE 270 @ 65.80

INVENTORY BALANCE

  

135 @ 53.80 = 7263

155@ 63.80 = 9889

20 @ 65.80 = 1316

18468

  

c) weighted average

  

MAR I

OP STOCK 240 @ 53.80 = 12912

MAR 5

PURCHASE 295 @58.80 = 17346 W A = (12912+17346)/(240+295) = 56.56

MARCH 9

SALES 400

INVENTORY BALANCE

135 @ 56.56 = 7636

MARCH 18

PURCHASE 155@ 63.80 = 9889

MARCH 25

PURCHASE 290@ 65.80 = 19082 WA = (7636+9889+19082)/(135+155+290)=63.12

INVENTORY BALANCE = 580 @ 63.12

MARCH 29

SALE 270

INVENTORY BALANCE = 580-270

= 310 @63.12 = 19567.2

D) SPECIAL

MAR I

OP STOCK 240 @ 53.80   

MAR 5

PURCHASE 295 @58.80

MARCH 9

SALES 135@ 53.80

265 @ 58.80

INVENTORY BALNCE

105 @ 53.80

30@ 58.80

MARCH 18

PURCHASE 155@ 63.80

MARCH 25

PURCHASE 290@ 65.80

INVENTORY BALANCE   

105 @ 53.80

30@ 58.80

155@ 63.80

  290@ 65.80

MARCH 29

SALE 115 @63.80

155 @ 65.80

INVENTORY BALANCE

  

  

105 @ 53.80 = 5649

30@ 58.80 = 1764

40@ 63.80 = 2552

135@ 65.80 = 8883

18848